A resistive position transducer with a resistance of 5000 and a shaft stroke of 5.0 in. is used in the arrangement of Fig. (4). Potentiometer R 3 R 4 is also 5000, and V T = 5.0 V. The initial position to be used as a reference point is such that R 1 =R 2 (i.e.. the shaft is at midstroke). At the start of the test, potentiometer R 3 R 4 is adjusted so that the bridge is balanced (V E =0). Assuming that the object being monitored will move a maximum distance of 0.5 in. toward A, what will the new value of V E be


1
Expert's answer
2021-01-25T13:58:30-0500

—If the wiper moves 0.5 in. toward A from midstroke, it will be 3.0 in. from B.

"R_2 = \large\frac{3.0 in}{5.0in}" * "5000\varOmega = 3000\varOmega"


"V_E = V_{R2} = V_{E4} - \large\frac{R_2}{R_1+R_2}" "V_T - \large\frac{R_4}{R_3+R_4}" =


"= \large(\frac{3000\varOmega}{5000\varOmega})(5V) - \large(\frac{2500\varOmega}{5000\varOmega})(5V) = 0.5V"

the new value of VE is 0.5V